OpenStudy (anonymous):

I need help with differentiating the function: f(x)=ln(sin^2x)

OpenStudy (anonymous):

\[f(x)=\ln (\sin ^2x)\]

OpenStudy (anonymous):

Start off by letting y = f(x) and rearranging to \[y= 2\ln{\sin(x)}\] Now, let \[u = \sin{x}\] \[\frac{du}{dx} = \cos{x}\] And y can now be written as: \[y = 2\ln{u}\] \[\frac{dy}{du} = \frac{2}{u}\] And by the chain rule: \[\frac{dy}{dx} = \frac{dy}{du}*\frac{du}{dx}\] So \[\frac{dy}{dx} = \frac{2}{u} * \cos{x}\] Replacing u with sin(x): \[\frac{dy}{dx} = \frac{2}{\sin{x}} * \cos{x}\] And tidying up: \[\frac{dy}{dx} = 2\cot{x}\]

OpenStudy (anonymous):

Thank you! I couldn't figure this prob out all the way :)